OpenStudy (crashonce):

Find a cubic equation whose roots are the squares of those of x^3+px+q=0. State reasons and working

ganeshie8 (ganeshie8):

Do you know how to find a polynomial whose roots are negative of the given polynomial ?

OpenStudy (crashonce):

not sure what you mean @ganeshie8

ganeshie8 (ganeshie8):

If r is a root if the given polynomial, then you should find a polynomial that has a root as -r

OpenStudy (crashonce):

so say thenormal poly was p(x) = (x-a)S(x)+R the new poly would be p(x) = (x+a)S(x)+r

OpenStudy (fwizbang):

Assume the solutions are a,b, and c, So that \[(x-a)(x-b)(x-c)=0\] The equation you're looking for is \[(x-a^2)(x-b^2)(x-c^2)=0\] The solution of the original depressed cubic is known so you can look up the relation between a,b,c, and p,q. See https://en.wikipedia.org/wiki/Cubic_function Once you have the three solutions, plug them into the 2nd equation and multiply it out.

ganeshie8 (ganeshie8):

Nope. If r is a root of p(x), the you can factor p(x) as : p(x) = (x-r)S(x)

OpenStudy (crashonce):

yo @fwizbang i dont realy get wat ur tryna point out. could you explain differently> thanks

ganeshie8 (ganeshie8):

Yes. So let's first focus on funding a polynomial whose roots are negative of the given polynomial

OpenStudy (crashonce):

how would we do that?

OpenStudy (fwizbang):

Every cubic polynomial has three roots, call them a,b, and c. Since the cubic is zero at each of the roots, we can write \[x^3-px+q = (x-a)(x-b)(x-c)=0\] If you write out the right hand side, you get \[a+b+c =0\]\[p=ab+bc+ac\] and \[q=-abc\] What happens to p and q if you change a,b, and c to -a,-b, and -c?

ganeshie8 (ganeshie8):

If r is a root of p(x), then what's the value of p(x) at x= r ?

OpenStudy (crashonce):

p is same so p is still ab+bc+ac and q becomes positive?

OpenStudy (pawanyadav):

After solving I got this equation.. x3+ 2p(x2) +p2x+q2=0 Is it correct?

OpenStudy (crashonce):

@Pawanyadav your answer is correct except it is -q^2 not +q^2 how did you do it?

OpenStudy (pawanyadav):

Oh sorry writing mistake

OpenStudy (crashonce):

how'd u get it

OpenStudy (pawanyadav):

(x-a2)(x-b2)(x-c2)=0 fwizbang posted this above

OpenStudy (crashonce):

then?

OpenStudy (fwizbang):

Do you have the roots(a,b,c). If you do, just multiply it out....

OpenStudy (pawanyadav):

Just simplify it And u will get.. x3-x2(a2+b2+c2)+x(a2b2+b2c2+c2a2)-a2b2c2=0

OpenStudy (crashonce):

where did you get the x(a2b2... term

OpenStudy (crashonce):

when i expanded i got: x3-x2(a2+B2+C2)+x(a2b2) did i do sth wrong

OpenStudy (pawanyadav):

And from fwizbang equation a+b+c=0 ab+bc+ca=p abc=-q So a2+b2+c2=-2p And (a2b2+........+c2a2)=0 Just substitute this in above equation.

OpenStudy (crashonce):

oh! i getit

OpenStudy (pawanyadav):

Try again....

OpenStudy (crashonce):

thanks!

OpenStudy (pawanyadav):

Thanks u got it